# IncogniText: Privacy-enhancing Conditional Text Anonymization via LLM-based Private Attribute Randomization

### v1

@@ Abstract
In this work, we address the problem of text anonymization where the goal is to prevent adversaries from correctly inferring private attributes of the author, while keeping the text utility, i.e., meaning and semantics. We propose IncogniText, a technique that anonymizes the text to mislead a potential adversary into predicting a wrong private attribute value. Our empirical evaluation shows a reduction of private attribute leakage by more than $90\%$. Finally, we demonstrate the maturity of IncogniText for real-world applications by distilling its anonymization capability into a set of LoRA parameters associated with an on-device model. 11footnotetext: Equal contribution, alphabetical order
[…]
envname-P envname#1
IncogniText: Privacy-enhancing Conditional Text Anonymization via LLM-based Private Attribute Randomization
Ahmed Frikha*  Nassim Walha* Krishna Kanth NakkaRicardo Mendes  Xue Jiang Xuebing Zhou   Huawei Munich Research Center ahmed.frikha1@huawei.com
@@ 1 Introduction
Large Language Models (LLMs), e.g., GPT-4 Achiam et al. (2023), are gradually becoming ubiquitous and part of many applications in different sectors, e.g., healthcare Liu et al. (2024) and law Sun (2023), where they act as assistants to the users. Despite their various benefits Noy and Zhang (2023), the power of LLMs can be misused for harmful purposes, e.g., cybersecurity Xu et al. (2024) and privacy Neel and Chang (2023) attacks, and profiling Brewster . For instance, LLMs were found to be able to predict various private attributes, e.g., age, gender, income, occupation, about the text author Staab et al. (2023). Hereby, they achieve a performance close to that of humans with internet access, while incurring negligible costs and time. Such private attributes are quasi-identifiers and their combination can substantially increase the likelihood of re-identification Sweeney (2000), i.e., revealing the text author identity. This suggests that human-written text data could in some cases be considered as personal data, which is defined as "any information relating to an identified or identifiable natural person" in GDPR European Parliament and Council of the European Union . Hence, human-written text might potentially require further analysis and protection measures to comply with such privacy regulations.
@@
Prior works proposed word-level approaches to mitigate text privacy leakage Albanese et al. (2023); Li et al. (2023). However, lexical changes do not change the syntactic features which were found to be sufficient for authorship attribution Tschuggnall and Specht (2014). Another line of work leverages differential privacy techniques to re-write the text in a privacy-preserving way Weggenmann et al. (2022); Igamberdiev and Habernal (2023), however, with high utility loss. Moreover, while most prior works and current state-of-the-art text anonymization industry solutions succeed in identifying and anonymizing regular separable text portions, e.g., PII, they fail in cases where intricate reasoning involving context and external knowledge is required to prevent privacy leakage Pilán et al. (2022). In light of this and given that most people do not know how to minimize the leakage of their private attributes, methods that effectively mitigate this threat are urgently needed.
@@
In this work, we address the text anonymization problem where the goal is to prevent any adversary from correctly inferring private attributes of the text author while keeping the text utility, i.e., meaning and semantics. This problem is a prototype for a practical use case where data can reveal quasi-identifiers about the text author, e.g., online services (ChatGPT) and anonymous social media platforms (Reddit). Our contribution is threefold: First, we propose a novel text anonymization method that substantially increases its protection against attribute inference attacks. Second, we demonstrate the effectiveness of our method by conducting an empirical evaluation with different LLMs and on 2 datasets. Here, we also show that our method achieves higher privacy protection compared to two concurrent works Dou et al. (2023); Staab et al. (2024). Finally, we demonstrate the maturity of our method for real-world applications by distilling its anonymization capability into a set of LoRA parameters Hu et al. (2022) that can be added to a small on-device model on consumer products.
@@ 2 Method
We propose IncogniText, an approach to leverage an LLM to protect the original text against attribute inference, while maintaining its utility, i.e., meaning and semantics, hence achieving a better privacy-utility trade-off. Given a specific attribute $a$, e.g., age, our method protects the original text $x_{orig}$ against the inference of the author’s true value $a_{true}$ of the private attribute, e.g., age: 30, by re-writing it in a way that misleads a potential privacy attacker into predicting a wrong target value $a_{target}$, e.g., age: 45. See Fig. 1 for an illustrative example.
We use an anonymization model $M_{anon}$ to re-write the original text $x_{orig}$ using a target attribute value $a_{target}$, the true attribute value $a_{true}$, and the template $T_{anon}$ with anonymization demonstrations, yielding $x_{anon}=M_{anon}(x_{orig},a_{target},a_{true},T_{anon})$. Hereby, the target value can either be chosen by the user or randomly sampled from a pre-defined set of values for the attribute considered. We additionally inform the anonymizer of the true attribute value $a_{true}$ to achieve an anonymized text $x_{anon}$ particularly tailored to hiding that value. The true attribute value could either be read from the text author’s device, e.g., local on-device profile or personal knowledge graph, or input by the author. Nevertheless, IncogniText achieves very effective anonymization even without the usage of the true attribute value $a_{true}$ as demonstrated by our experiments (Section 3).
To validate the effectiveness of the anonymized text $x_{anon}$ against attribute inference, we use a simulated adversary model $M_{adv}$ that tries to predict the author’s attribute value $a_{true}$. If the prediction is correct, additional rounds of anonymization are conducted with the anonymization model $M_{anon}$ until the adversary model $M_{adv}$ is fooled or a maximum iteration number is reached. This ensures that we perform as few re-writing iterations as necessary, hence maintaining as much utility as possible, i.e., the original text is changed as little as possible. Note that the same model can be used as $M_{anon}$ and $M_{adv}$ with different prompt templates $T_{anon}$ and $T_{adv}$ respectively (see Appendix). This might be especially suitable for on-device anonymization cases with limited memory and compute. Note that applying IncogniText to multiple attributes is easily achieved by merging the attribute-specific parts of the anonymization templates. For cases where the text author wants to share a subset or none of the private attributes, they can flexibly choose which attributes to anonymize, if any.
In addition to its usage for early stopping of the iterative anonymization, the adversary model $M_{adv}$ can also be used to inform the anonymization by sharing its reasoning $x_{AdvReasoning}$ for the correct prediction of the true attribute value $a_{true}$. In this case, the reasoning text $x_{AdvReasoning}$ is fed as an additional input to the anonymization model $M_{anon}$. This feature was also proposed in the concurrent work Staab et al. (2024) and we evaluate this variant of IncogniText in our experimental study. We highlight the main differences between this concurrent work and our approach. First, we condition the anonymization model $M_{anon}$ on a target attribute value $a_{target}$. We believe that misleading a potential attacker into predicting a wrong private attribute value by inserting new hints is more effective than removing or abstracting hints to the original value present in the original text. Furthermore, we condition the anonymization model $M_{anon}$ of the true attribute value $a_{true}$ to increase the quality of the anonymization. Finally, we leverage the adversary model $M_{adv}$ as an early stopping method to prevent unnecessary utility loss or the deterioration of the anonymization quality, i.e., further anonymization iterations can in some cases lead to a decrease in privacy as observed in the experiments in Staab et al. (2024). Our empirical evaluation and ablation study demonstrate the effectiveness of these contributions.
@@
Figure 1: IncogniText example: The true user attribute value (middle income) is obfuscated by replacing it with a wrong target value (low income) with minimal text changes.
@@ 3 Experimental evaluation
We first evaluate our approach on the dataset of $525$ human-verified synthetic conversations proposed by Staab et al. (2023). The dataset includes 8 different private attributes: age, gender, occupation, education, income level, relationship status, and the country and city where the author was born and currently lives in. We compare to anonymization baseline approaches including the Azure Language Service (ALS) Aahill (2023) and the two concurrent works, Dou-SD Dou et al. (2023) and Feedback-guided Adversarial Anonymization (FgAA) Staab et al. (2024). We evaluate the privacy of the anonymized texts using the SOTA attribute inference attack method Staab et al. (2023) which leverages pre-trained LLMs to predict the author attributes based on the text. We assess the utility of the anonymized texts using the traditional ROUGE score Lin (2004) and the LLM-based utility evaluation with the utility judge template $T_{utility}$ proposed in Staab et al. (2024). The latter computes the mean of scores for meaning, readability, and hallucinations given by the evaluation model. More details about the experimental setting including the prompt templates can be found in the Appendix E.
@@
Synthetic Reddit-based dataset Staab et al. (2023)
@@
ALS Aahill (2023) 55 96 64
Dou-SD Dou et al. (2023) 47 64 78
FgAA Staab et al. (2024) 26 68 86
FgAA Staab et al. (2024) 43.2 87.9 98.8
@@
Real self-disclosure dataset Dou et al. (2023)
@@
Table 1: Attribute-averaged results (%) of attacker attribute inference accuracy (Privacy), ROUGE-score, and LLM judge score (Utility). Results denoted by are reported from Staab et al. (2024) where the anonymized texts were evaluated by GPT-4. Results denoted by are our reproductions where the anonymized texts were evaluated with Phi-3-small. For FgAA, we use the best anonymizer model in our experiments (Phi-3 small).
@@
Table 1 presents our main results. We find that IncogniText achieves the highest privacy protection, i.e., lowest attacker inference accuracy, with a tremendous improvement of ca. $19\%$ compared to the strongest baseline. Note that FgAA uses a stronger anonymizer model (GPT-4) suggesting that the improvement might be bigger if we would use the same model with our method. Most importantly, we find that IncogniText substantially reduces the amount of attribute value correctly predicted by the attacker by ca. $90\%$, namely from $71.2\%$ to $7.2\%$. Moreover, our approach achieves high privacy protection across different model sizes and architectures, i.e., Llama 3 Meta (2024) and Phi-3 Abdin et al. (2024), demonstrating that it is model-agnostic. While the IncogniText-anonymized texts yield a high utility, we find that our reproduction of FgAA achieves higher utility scores. This is explained by the lower meaning and hallucination scores (the more the model hallucinates, the lower its hallucination score, see Appendix) assigned to IncogniText-anonymized texts by the LLM-based utility judge which considers the inserted cues to mislead the attacker as hallucinations. We argue that these changes are desired by the text author and that they are required to successfully fool the attacker into predicting a wrong attribute value. Finally, we find that IncogniText is significantly faster than the baseline, effectively requiring less anonymization steps (Fig. 2).
We also validate our approach on a dataset proposed in the concurrent work Dou et al. (2023) which contains real posts and comments from Reddit with annotated text-span self-disclosures. We choose a subset of 196 examples that we preprocess (see Appendix for details). Likewise, IncogniText significantly outperforms the strongest baseline FgAA on this dataset, reducing the adversarial accuracy by ca. 82%, namely from 73% to 12.8%.
@@
We investigate different variants of IncogniText to gain more insights into the importance of its components and present the results in Table 2. First, we observe that conditioning the anonymization on a target attribute value is crucial for achieving high privacy protection. Besides, we find that performing early stopping (ES) with the adversary model improves both privacy and utility, since it ensures that no further anonymization steps are conducted that might deteriorate utility or privacy. Moreover, our results suggest that conditioning the anonymizer (Anon) on the attribute ground truth (GT) value is more important than conditioning it on the adversarial reasoning and inference (Inf) for achieving higher privacy. In contrast, conditioning the adversary (Adv) on the GT deteriorates all metrics. We hypothesize that the adversary identifies fewer cues about the author in the text when it has access to GT. Ablation results with other models as anonymizers and Phi-3-small as evaluation model (see appendix) also showcase that conditioning on a target value is the main factor for decreasing privacy leakage. However, their results show no clear trend for the effect of conditioning the anonymizer and the adversary on any other information (GT or Inf). Results that were reported in Table 1 correspond to the $5^{th}$ experiment from table 2.
@@
Table 2: Attribute-averaged results (%) of the ablation study with Phi-3-small as anonymization and evaluation model. Examined components: 1) using the target wrong attribute value $a_{target}$ (Target), 2) conditioning the anonymizer (Anon) on the inference reasoning of the adversary (Inf), on the ground truth (GT) attribute value, or both, 3) whether to condition the adversary model (Adv) on GT, 4) using the adversary to perform early stopping (ES), i.e., stop the iterative anonymization once it predicts the attribute value incorrectly.
Finally, we investigate whether IncogniText can achieve a high privacy protection as part of an on-device anonymization solution. For this, we distill the IncogniText anonymization capabilities of the best anonymizer model (Phi-3-small) into a dedicated set of LoRA Hu et al. (2022) parameters associated with a small Qwen2-1.5B model Bai et al. (2023) that could be run on-device. We perform the instruction-finetuning Wei et al. (2021) using additional synthetic conversations released by Staab et al. (2023) that are different than the 525 examples used for testing. The additional examples were not included in the officially released set due to quality issues, e.g., wrong formatting, hallucinations, or absence of hints to the private attributes. We filter and post-process this set of data to solve the issues yielding 664 new examples to which we apply IncogniText to create input-output pairs that we use for finetuning and validation. Post-processing details can be found in the Appendix. We finetune the anonymizer to perform the $4^{th}$ experiment in Table 2 (Anonymizer conditioned on Target and GT). We only fine-tune the anonymizer model and use the pretrained version of Qwen2-1.5B for the adversary. The results (Table 3) show a substantial privacy improvement on-device, effectively reducing the private attribute leakage by more than 50%, from 40.8% to 18.1%, while maintaining utility scores comparable to larger models.

### v2

# envname-P envname#1

@@ Abstract
In this work, we address the problem of text anonymization where the goal is to prevent adversaries from correctly inferring private attributes of the author, while keeping the text utility, i.e., meaning and semantics. We propose IncogniText, a technique that anonymizes the text to mislead a potential adversary into predicting a wrong private attribute value. Our empirical evaluation shows a reduction of private attribute leakage by more than $90\%$ across 8 different private attributes. Finally, we demonstrate the maturity of IncogniText for real-world applications by distilling its anonymization capability into a set of LoRA parameters associated with an on-device model. Our results show the possibility of reducing privacy leakage by more than half with limited impact on utility. 11footnotetext: Equal contribution, alphabetical order
@@ 1 Introduction
Large Language Models (LLMs), e.g., GPT-4 [3], are gradually becoming ubiquitous and part of many applications in different sectors, e.g., healthcare [15] and law [24], where they act as assistants to the users. Despite their various benefits [18], the power of LLMs can be misused for harmful purposes, e.g., attacks on cybersecurity [29] and privacy [17], as well as profiling [6]. For instance, LLMs were found to be able to predict various private attributes, e.g., age, gender, income, occupation, about the text author [22]. Hereby, they achieve a performance close to that of humans with internet access, while incurring negligible costs and time. Such private attributes are quasi-identifiers and their combination can substantially increase the likelihood of re-identification [25], i.e., revealing the text author identity. This suggests that human-written text data could in some cases be considered as personal data, which is defined as "any information relating to an identified or identifiable natural person" in GDPR [10]. Hence, human-written text might potentially require further analysis and protection measures to comply with such privacy regulations.
@@
Prior works proposed word-level approaches to mitigate text privacy leakage [4, 13]. However, lexical changes do not change the syntactic features which were found to be sufficient for authorship attribution [26]. Another line of work leverages differential privacy techniques to re-write the text in a privacy-preserving way [27, 12], however, with high utility loss. Moreover, while most prior works and current state-of-the-art text anonymization industry solutions [20] succeed in identifying and anonymizing regular separable text portions, e.g., PII, they fail in cases where intricate reasoning involving context and external knowledge is required to prevent privacy leakage [20]. In light of this and given that most people do not know how to minimize the leakage of their private attributes, methods that effectively mitigate this threat are urgently needed.
@@
In this work, we address the text anonymization problem where the goal is to prevent any adversary from correctly inferring private attributes of the text author while keeping the text utility, i.e., meaning and semantics. This problem is a prototype for a practical use case where data can reveal quasi-identifiers about the text author, e.g., user interaction with online services (ChatGPT) and user content in anonymous social media platforms (Reddit). Our contribution is threefold: First, we propose a novel text anonymization method that substantially increases its protection against attribute inference attacks. Second, we demonstrate the effectiveness of our method by conducting an empirical evaluation with different LLMs and on 2 datasets. Here, we also show that our method achieves higher privacy protection compared to two concurrent works [7, 23]. Finally, we demonstrate the maturity of our method for real-world applications by distilling its anonymization capability into a set of LoRA parameters [11] that can be added to a small on-device model on consumer products.
@@ 2 Method
The threat model we consider in the present paper is as follows: Under an anonymous pseudonym, the user publishes some text on a platform, e.g., Reddit or Twitter, or sends it as part of a request to an online-service, e.g., ChatGPT. We assume that the adversary is anyone with access to the published text, which includes users, platform admins, and general public, in the case of an online platform, and the service provider in the case of an online-service. Based on the text, the adversary attempts to infer information from the user, particularly their private attributes (profiling) such as, gender, age, location, nationality, etc. We further assume that only this particular text is accessible. This simplification is made to focus the scope of the present work on text anonymization, while leaving out privacy issues stemming from metadata, e.g., IP or fingerprinting. However, we note that our proposed method can be applied to multiple pieces of text, e.g., via a simple concatenation of the text pieces.
@@
Figure 1: Overview of the IncogniText anonymization approach. In this example, the true user attribute value $a_{true}$ (middle income) is obfuscated via the IncogniText anonymization conditioned on a wrong target value $a_{target}$ (low income) with minimal text changes.
Our approach, IncogniText, leverages an LLM to protect the original text against attribute inference, while maintaining its utility, i.e., meaning and semantics, hence achieving a better privacy-utility trade-off. Given a specific attribute $a$, e.g., age, our method protects the original text $x^{orig}$ against the inference of the author’s true value $a_{true}$ of the private attribute, e.g., age: 30, by re-writing it to an anonymized version $x^{anon}$ that misleads a potential privacy attacker into predicting a predefined wrong target value $a_{target}$, e.g., age: 45. See Fig. 1 for an illustrative example of another private attribute (income level). IncogniText is composed of two stages which are executed iteratively (Fig. 1), in a way that mirrors an adversarial training paradigm. Algorithm $1$ describes our approach.
In the first stage, we use an adversarial model $M_{adv}$ to predict the author’s attribute value $a_{true}$ along with a reasoning $R$ that supports its inference $I$, i.e., the predicted value. This is done by using the adversarial inference template $T_{adv}$ (see Appendix C) and the user text. While in the first anonymization iteration, the user text used is the original text $x^{orig}$ written by the user, in further iterations, the output of the second stage is used instead, i.e., the anonymized version of the user text $x^{anon}_{i-1}$ of the iteration $i-1$ is fed to the adversarial model $M_{adv}$. It is important to note that the adversarial model $M_{adv}$ is different from the model $M_{attacker}$ that might be used by a potential attacker. The objective of using the adversarial model is to locally mimic the attack that a potential attacker might conduct and leverage the output of the attack to enhance the anonymization step applied in stage 2. We leverage the adversarial model $M_{adv}$ as a proxy of the potential attacker model $M_{attacker}$. In our experiments (Section 3), we use different models for the adversarial and attacker models, $M_{adv}$ and $M_{attacker}$, respectively.
In the second stage, we employ an anonymization model $M_{anon}$ to further anonymize the user text $x^{anon}_{i-1}$ of the iteration $i-1$ by using the anonymization template $T_{anon}$, a target attribute value $a_{target}$, and reasoning $R$ and inference $I$ generated by the adversarial model, yielding a new anonymized version of the user text $x^{anon}_{i}$. Our choice to leverage a target attribute value $a_{target}$ is based on the intuition that misleading a potential attacker into predicting a wrong private attribute value by inserting new hints is more effective than removing or abstracting hints to the original value present in the original text. Hereby, the target value $a_{target}$ can either be chosen by the user or randomly sampled from a pre-defined set of values for the attribute considered. Besides, we use the reasoning $R$ and inference $I$ generated by the adversarial model $M_{adv}$ to inform the anonymization process, in a way that mimicks adversarial training. Furthermore, we considered further variants of Incognitext where we additionally inform the anonymization model $M_{anon}$ of the true attribute value $a_{true}$ (not shown in Fig. 1) to achieve anonymized texts $x^{anon}_{i}$ particularly tailored to hiding that value. In practice, the true attribute value could either be read from the text author’s device, e.g., local on-device profile or personal knowledge graph, or input by the author manually. Nevertheless, IncogniText achieves very effective anonymization even without the usage of the true attribute value $a_{true}$ as demonstrated by our experiments (Section 3).
Algorithm 1 IncogniText Anonymization
0:  $M_{anon},M_{adv}$: Anonymization and adversary models
0:  $T_{anon},T_{adv}$: Anonymization and adversary prompt templates
0:  $x^{orig}$: Original user text
0:  $a_{target}$: Target attribute value
0:  $a_{true}$: True attribute value of the user (Optional)
0:  $n$: Maximum number of anonymization iterations
1:  $x^{anon}_{0}=x^{orig}$ // Initialize the anonymized text to the original text
2:  for $i=1..n$ do
3:     $I,R=M_{adv}(x^{anon}_{i-1},T_{adv})$ // Get inference I and reasoning R via adversarial evaluation
4:     if $I=a_{true}$ then
5:        $x^{anon}_{i}=M_{anon}(x^{anon}_{i-1},T_{anon},a_{target},a_{true},I,R)$ // Perform an anonymization iteration
6:     else
7:        break // Early stopping
8:     end if
9:  end for
10:  Return Anonymized user text $x^{anon}_{i-1}$
The iterative application of both stages is executed as long as the inference $I$ predicted by the adversary model $M_{adv}$ matches the user true attribute value $a_{true}$, i.e., the adversarial inference is used as early stopping criterion, or a maximum iteration number is reached. This ensures that we perform as few re-writing iterations as necessary, hence maintaining as much utility as possible, i.e., the original text is changed as little as possible.
Note that the same model $M$ can be used as $M_{anon}$ and $M_{adv}$ with different prompt templates $T_{anon}$ and $T_{adv}$ respectively. This is especially suitable for on-device anonymization cases with limited memory and compute. Note that applying IncogniText to multiple attributes can be easily achieved by merging the attribute-specific parts of the anonymization templates (see Appendix for details). For cases where the text author wants to share a subset or none of the private attributes, they can flexibly choose which attributes to anonymize, if any.
The main difference to prior approaches focusing on direct leakage of PII or private attributes that are explicitly mentioned in the text [1, 7] is that IncogniText protects also against indirect leakage. By leveraging the reasoning abilities of the LLM, our approach detects any forms of indirect leakage attribute/PII value, e.g., hints or cues that might be combined with external knowledge to infer the private information, identifies the text spans responsible for the correct PII/attribute inference, and then modifies these text spans to mislead a potential attacker into predicting the target value. Most importantly, these three operations are performed end-to-end implicitly by our LLM-based approach, without training for these steps or separating them conceptually. IncogniText adopts a rewriting approach instead of a detect-and-replace approach. Our method can be viewed as performing replacement of the attribute values in the latent attribute space instead of doing so in the text representation space.
@@ 3 Experimental evaluation
The datasets we used to evaluate our approach include a dataset of $525$ human-verified synthetic conversations proposed by staab2023beyond [22] and a dataset proposed in the concurrent work dou2023reducing [7] which contains real posts and comments from Reddit with annotated text-span self-disclosures. The former dataset contains 8 different private attributes: age, gender, occupation, education, income level, relationship status, and the country and city where the author was born and currently lives in. For the second dataset, we consider the following attributes: gender, relationship status, age, education, and occupation. We keep only samples where the author discloses information about their own private attributes and not about someone else. Furthermore, we label the samples with the real private attribute values instead of text spans, yielding a set of 196 examples.
The baseline methods we compared to in our evaluation include the Azure Language Service (ALS) [1], the differentially private DP-BART-PR+ [12] and the two concurrent works, Dou-SD [7] and Feedback-guided Adversarial Anonymization (FgAA) [23]. DP-BART-PR+ is a variant of the differentially private method DP-BART that uses the encoder-decoder model BART [bart] to rewrite arbitrary input text sequences while providing local differential privacy guarantees for the output text. To provide these guarantees and preserve text utility, the encoder output neurons are first pruned, then the resulting encoder output is clipped by value, and finally, noise is added to the clipped outputs before being sent to the decoder. In our experiments, we use DP-BART-PR+ with privacy budgets $\varepsilon=1000$ and $\varepsilon=2500$ following the value ranges used by igamberdiev2023dp [12]. Dou-SD is an approach that uses an LLM to detect voluntary self-disclosure of private attributes and personal information, then finetunes a model to replace the detected text spans with anonymized versions. FgAA leverages a pretrained LLM to perform the anonymization in an adversarial manner where an adversarial proxy is used to inform the anonymization process, similarly to our method. In the following, we highlight the main differences between this concurrent work and our approach. First, we condition the anonymization model $M_{anon}$ on a target attribute value $a_{target}$. We believe that misleading a potential attacker into predicting a wrong private attribute value by inserting new hints is more effective than removing or abstracting hints to the original value present in the original text. Furthermore, we condition the anonymization model $M_{anon}$ of the true attribute value $a_{true}$ to increase the quality of the anonymization. Finally, we leverage the adversary model $M_{adv}$ as an early stopping method to prevent unnecessary utility loss or the deterioration of the anonymization quality, i.e., further anonymization iterations can in some cases lead to a decrease in privacy as observed in the experiments in [23]. Our empirical evaluation and ablation study demonstrate the effectiveness of these contributions.
The evaluation metrics we used can be categorizes in privacy and utility evaluation methods. For the privacy evaluation of the anonymized texts, we used an attribute inference attack method which leverages pre-trained LLMs to predict the author attributes based on the text [22]. The privacy metric used is the prediction accuracy of the attacker that uses this method. We use this LLM-based attribute inference attack because it is the SOTA attack for this category and it is the most suitable attack for free-form texts that do not explicitly mention either the attribute of the user or its values. In fact, the attributes and their value are mostly implicitly included in the text or require a complex reasoning combined with external knowledge to be inferred. We argue that such free-form texts are more representative of real-world scenarios and that this privacy evaluation method is what a potential attacker would realistically use in the light of the current literature. Other attacks require training specific attacker models, which requires labeled datasets for each attribute including examples from all classes (the attribute values) [21, 19, 9]. In our privacy evaluation, we do not assume that the attacker model is known and use different adversarial and attacker models, $M_{adv}$ and $M_{attacker}$, respectively. We use the strongest LLM in private attribute inference attacks, based on our experiments, as the attacker model $M_{attacker}$. In our experiments, we set the adversarial model $M_{adv}$ to always be the same as the anonymization model $M_{anon}$, which is suitable for low compute environments, e.g., on-device anonymization. Regarding the utility evaluation of the anonymized texts, we use the traditional ROUGE score [14] and the LLM-based utility evaluation with the utility judge template $T_{utility}$ proposed by staab2024large [23]. The latter computes the mean of scores for meaning, readability, and hallucinations given by the evaluation model. More details about the experimental setting including the prompt templates can be found in the Appendix C.
@@
Synthetic Reddit-based dataset [22]
@@
ALS [1] 55 96 64
DP-BART-PR+1000 [12] 27.81 7.05 39.07
DP-BART-PR+2500 [12] 62.48 93.36 91.87
Dou-SD [7] 47 64 78
FgAA [23] 26 68 86
FgAA [23] 43.2 87.9 98.8
@@
Real self-disclosure dataset [7]
@@
Table 1: Attribute-averaged results (%) of attacker attribute inference accuracy (Privacy), ROUGE-score, and LLM judge score (Utility). Results denoted by are reported from staab2024large [23] where the anonymized texts were evaluated by GPT-4. Results denoted by are our reproductions where the anonymized texts were evaluated with Phi-3-small. We also include results for DP-BART-PR+ with privacy budgets $\varepsilon=1000$ and $\varepsilon=2500$.
For FgAA, we use the best anonymization model in our experiments (Phi-3 small).
@@
Our main results are presented in Table 1. We find that IncogniText achieves the highest privacy protection, i.e., lowest attacker inference accuracy, with a tremendous improvement of ca. $19\%$ compared to the strongest baseline. Note that FgAA uses a stronger anonymization model (GPT-4) suggesting that the improvement might be bigger if we would use the same model with our method. Most importantly, we find that IncogniText substantially reduces the amount of attribute value correctly predicted by the attacker by ca. $90\%$, namely from $71.2\%$ to $7.2\%$. Moreover, our approach achieves high privacy protection across different model sizes and architectures, i.e., Llama 3 [16] and Phi-3 [2], demonstrating that it is model-agnostic. While the IncogniText-anonymized texts yield a high utility, we find that our reproduction of FgAA achieves higher utility scores. This is explained by the lower meaning and hallucination scores (the more the model hallucinates, the lower its hallucination score, see Appendix) assigned to IncogniText-anonymized texts by the LLM-based utility judge which considers the inserted cues to mislead the attacker as hallucinations. We argue that these changes are desired by the text author and that they are required to successfully fool the attacker into predicting a wrong attribute value. Although DP-BART-PR+ ($\epsilon=2500$) achieves a higher ROUGE score than IncogniText, it comes at the cost of providing almost no privacy protection. In fact, the very high privacy budget of 2500 results in almost no changes to the original text. We also note that IncogniText significantly outperforms the strongest baseline FgAA on the second real Reddit comments dataset [7], reducing the adversarial accuracy by ca. 82%, namely from 73% to 12.8%.
Figure 2: Private attribute inference accuracy (%) by attribute for unprotected text and different anonymization methods. The first four bars for each attribute represent accuracy values reported in [23] and are evaluated using GPT-4 as the privacy evaluation model. The remaining five bars are evaluations from our experiments, performed using Phi-3-small as the privacy evaluation model.
Figure 2 illustrates our private attribute inference accuracy by attribute results. Based on the results of our implementation, which are evaluated using Phi-3-small, we observe that overall IncogniText yields a higher privacy protection, i.e., a lower private attribute inference accuracy, than all other methods across all attributes. One exception is observed for the Occupation (4%), where DP-BART-PR+ ($\epsilon=1000$) achieves a private attribute inference accuracy of 0%. However, a closer look at the rewritten output of DP-BART-PR+ ($\epsilon=1000$) reveals a complete loss of the text semantics due to excessive noise added, and therefore a complete loss of utility. This is confirmed by the extremely low utility scores in table 1. Moreover, IncogniText reduces at least 90% of the original privacy leakage in all attributes, except for Income Level, where 86% of the original privacy leakage is reduced after rewriting. In contrast, FgAA prevents less than 50% of the original privacy leakage for the attributes Relationship status, Sex, and Income level, with only a reduction of only 8% in private attribute inference accuracy for the latter.
Finally, we find that IncogniText is significantly faster than the baseline, effectively requiring less anonymization steps (Fig. 3). In fact, more than 80% of the test samples are successfully anonymized after a single iteration using our method, while more than half of the samples require a second and possibly a third iteration in FgAA.
@@
Our investigation of different IncogniText variants was conducted to gain more insights into the importance of its components. We present its results in Table 2. First, we observe that conditioning the anonymization on a target attribute value is crucial for achieving high privacy protection. Besides, we find that performing early stopping (ES) with the adversary model improves both privacy and utility, since it ensures that no further anonymization steps are conducted that might deteriorate utility or privacy. Moreover, our results suggest that conditioning the anonymization model (Anon) on the attribute ground truth (GT) value is more important than conditioning it on the adversarial reasoning and inference (Inf) for achieving higher privacy. In contrast, conditioning the adversary (Adv) on the GT deteriorates all metrics. We hypothesize that the adversary identifies fewer cues about the author in the text when it has access to GT. Ablation results with other models as anonymization models and Phi-3-small as evaluation model (see appendix) also showcase that conditioning on a target value is the main factor for decreasing privacy leakage. However, their results show no clear trend for the effect of conditioning the anonymization model and the adversary on any other information (GT or Inf). Results that were reported in Table 1 correspond to the $5^{th}$ experiment from table 2.
@@
Table 2: Attribute-averaged results (%) of the ablation study with Phi-3-small as anonymization and evaluation model. Examined components: 1) using the target wrong attribute value $a_{target}$ (Target), 2) conditioning the anonymization model (Anon) on the inference reasoning of the adversary (Inf), on the ground truth (GT) attribute value, or both, 3) whether to condition the adversary model (Adv) on GT, 4) using the adversary to perform early stopping (ES), i.e., stop the iterative anonymization once it predicts the attribute value incorrectly.
On-device anonymization experiments were conducted to investigate whether IncogniText can achieve a high privacy protection as part of an on-device anonymization solution. For this, we distill the IncogniText anonymization capabilities of the best anonymization model (Phi-3-small) into a dedicated set of LoRA [11] parameters associated with a small Qwen2-1.5B model [5] that could be run on-device. We perform the instruction-finetuning [28] using additional synthetic conversations released by [22] that are different than the 525 examples used for testing. The additional examples were not included in the officially released set due to quality issues, e.g., wrong formatting, hallucinations, or absence of hints to the private attributes. We filter and post-process this set of data to solve the issues yielding 664 new examples to which we apply IncogniText to create input-output pairs that we use for finetuning and validation. Post-processing details can be found in the Appendix. We finetune the anonymization model to perform the $4^{th}$ experiment in Table 2 (anonymization model conditioned on Target and GT). We only fine-tune the anonymization model and use the pretrained version of Qwen2-1.5B for the adversary. The results (Table 3) show a substantial privacy improvement on-device, effectively reducing the private attribute leakage by more than 50%, from 40.8% to 18.1%, while maintaining utility scores comparable to larger models.
@@
In our experiments, we sampled the target attribute value $a_{target}$ to be different from the real attribute value $a_{true}$. However, an adversary observing multiple anonymized texts from the same user might accumulate sufficient information to infer the real attribute value (by its absence). Note that this is a different threat model than the one we consider in the present work. To tackle this, we could easily adapt IncogniText to sample the target value $a_{target}$ uniformly from a the set of all possible attribute values (including the true one). To provide differential privacy (DP) guarantees, the solution can be implemented as the Randomized Response DP-mechanism [8] where the true target value $a_{target}$ is sampled with a pre-defined probability. We leave such endeavour for future work.